# Not Everyone Wins with LLMs: Behavioral Patterns and Pedagogical Implications in AI-assisted Data Analysis

### v1

@@ 4.1. Participants
Figure 2. Distribution of self-rated technical expertise, LLM expertise, and communication expertise scores.
A total of 41 students completed HW0, with a slight attrition to 36 students by HW4. Students came from a wide range of fields, including finance, telecommunications, journalism, physics, music, and computer science. This diversity highlights both the promise and the challenge of designing AI supports that work for “everyone.”
We collected pre-survey data on different self-reported expertise, including implementation (e.g., Python proficiency), domain (e.g., data science expertise), toolkit (e.g., Colab/Jupyter Notebook, pandas, and altair, which are the specific environment and libraries used for our tasks), model (e.g., GenAI/LLM proficiency), and non-technical skills (e.g, communication abilities). We initially aimed to analyze whether these factors influence effective LLM use in different ways; however, in our class, some skill areas were moderately correlated (e.g., a Spearman correlation $\rho=0.52$ for Python and data science items, and $\rho=0.72$ between data science and tookit items), making it difficult to disentangle their independent contributions (we discuss more about future work in Section 8.2).
Therefore, we operationalize students’ technical expertise by grouping the survey answers to programming, data science, and toolkit expertise questions, such as “How many years of experience do you have in using the following tools? [Python]” and “How confident are you in your ability to do the following tasks on your own? [Exploratory data analysis]” (more details in Appendix C). Similarly, we merged highly correlated items into composite measures of LLM expertise (e.g., “How would you assess your level of expertise in the following areas? [GenAI/LLM]”) and communication expertise (e.g., “I am confident in my communication ability in general.”). Items that were highly skewed toward zero or not informative were excluded. The distribution of these expertise among students is shown in Figure 2, where students were classified as novices or experts using a median split of self-ratings.
@@ 4.2. Data Collection
We administered structured surveys at multiple points in the course (see Appendix B for full questions):
  • Pre-Survey (start of semester): Collected demographics, academic and professional backgrounds, Likert-scale questions on self-perceived expertise, interest, confidence in personal vs. GenAI ability in different fields, etc.
@@
  • Final Post-Survey (end of semester): Collected self-assessed expertise, interest, and confidence across domains (similar to pre-survey). Probed on similarities and differences across tasks, what they learned, and what they still wanted to learn.
   "event_id": "evt_67",
   "hw": "hw0",
   "studentid": [student id],
   "timestamp": 1741801351065,
   "time_since_started": 3642517,
   "event_category": "prompt",
   "event_type": "converse",
   "event_name": "enter_prompt",
   "event_content": "What are the prices for Honda
                     cars manufactured after the year 2015",
   "task": "t2_eda",
   "converse_chatid": "chat5",
   "turn": "3",
   "event_response": [LLM Output]
Figure 3. An example log event format.
@@
We deployed a custom instrumentation pipeline in Colab to capture both notebook interactions (code executions/errors), selected UI events (pressing explain-code/error buttons) and LLM use (prompts/responses). Students installed a Tampermonkey script that intercepted requests to Colab’s converse and generateCode endpoints and captured clickstream events. The script added a “Download History” button to the Colab toolbar and stored code cells, chat messages, event metadata, and responses to browser localStorage, which students exported as JSON files upon submission (or auto-save when the storage is full).
In total, we collected 16,315 logged events (see example in Figure 3) and 44 hours of screen recordings across five assignments. Logs captured a rich sequence of AI interactions, code edits, and executions, supporting both quantitative distributional analyses and qualitative episode-based coding.
4.3. Data Analysis
Throughout the paper, we conducted a mixture of quantitative and qualitative analyses to answer different research questions. Here, we only provide an overview of the analysis methods, and we detail the method used to answer each research question in its corresponding sections:
  • RQ1: To assess whether different expertise makes a significant difference in grades (Section 5), we conduct regression testing on student homework data and competency scores;
  • RQ2: To characterize students’ AI usage behavior in relation to their task completion sequences, competency, and grades (Section 6), we perform thematic analysis identifying key AI behavior types, use them to annotate student logs, and then quantitatively analyze the resulting distributions;
  • RQ3: To further identify skills essential for effective AI usage (Section 7), we qualitatively analyzed survey responses as well as the aforementioned student log behaviors, identifying skill challenges and learning opportunities across conceptual, procedural, dispositional, and metacognitive dimensions.
5. RQ1: Do LLMs close different expertise gaps?
Analysis method: Quantitative regression modeling.
We assess whether different expertise (Figure 2, details about expertise score calculation in Section 4.1) makes a significant difference in grades in HW1–HW4. We fit a linear mixed-effects regression model with grades (normalized scores) as the outcome and technical, LLM, and communication expertise scores as predictors, including random intercepts for participant and assignment (Bates et al., 2015).
@@ 5. RQ1: Do LLMs close different expertise gaps?
Table 2. Linear mixed-effects regression predicting homework performance from different expertise. Participant and assignment were used as random intercepts.
@@
Analysis results: Technical experts can use LLMs more effectively than others.
@@
As shown in (Table 2), only technical expertise significantly predicted higher grades ($\beta=6.09,p=.041$). LLM and communication expertise had no significant effects and did not predict grades. This suggests that even with equal access to the Colab Gemini assistant, students with stronger programming and data science backgrounds continued to outperform peers.
@@
Note that students in our class had a median of 4/5 for their communication expertise (see Figure 2), which means that the communication factor may not have a lot of variance to predict grades. Meanwhile, students’ median of LLM expertise is 3.67/5, but everyday use of LLMs or their self-rated LLM knowledge may not straightforwardly transfer to complex programming data science tasks. Additionally, the expertise scores here are all operationalized as self-reported survey items, and there are some known issues in miscalibration of self-perceived expertise (Kruger and Dunning, 1999). Future work should explore more rigorous ways to measure expertise and recruit a more diverse set of students (more in Section 8.2).
Logs and surveys showed that students engaged substantially with AI during coursework: 15% of their logged events were prompt-related, and homework post-surveys all confirmed that students spent non-trivial amounts of time with AI tools. Despite substantial AI usage, the technical expertise gap remained — technical experts seem to benefit more from LLM use than novices, which is consistent with some prior literature (Chen et al., 2024).
6. RQ2: How do differences in expertise translate to AI usage behaviors?
@@
The analysis in RQ1 showed that students’ self-perceived technical expertise continued to predict performance even when all students had access to the Colab Gemini assistant. Here, we dive into student log data to examine why: how students use AI during workflows, and how it relates to different levels of technical expertise?
6.1. Analysis method: Thematic analysis and LLM-automated annotation
6.1.1. Data filtering.
As mentioned in Section 3, we required students to use the built-in Gemini and avoid using other LLMs. We added the precaution of asking students to self-report their additional LLM model use, if any. Despite our best efforts, we still noticed students using external AI tools, which tend to result in reasonable grades but no or few user logs captured, or logs with abnormal behaviors, e.g., copy-pasted large blocks of code with unfounded originality could not be found.
To ensure data quality for log analysis, we excluded these cases and retained logs from 28 students (16 experts [^2] The expert vs. novice for the rest of the paper are with respect to technical expertise., 12 novices) across 36 submissions (20 from experts, 16 from novices) across HW0, 1, and 2 (7,315 logged events). Specifically, we filtered out students who self-reported other tool use like ChatGPT or displayed other tool use in recording (if available), as well as those who disengaged or spent too much time working on tasks without AI, reflected as log length being less than the median. We also excluded logs from HW3-4 to enable meaningful comparison with HW0, where most students did not attempt Task 3 and 4.
Figure 4. (A) Average frequency of AI use codes per student (refer to Table 3 for code definitions), aggregated across all steps and clustered by AI usage code types, including improve, code, evaluate, explain, and prompt (quality). (B) Example annotated log snippet. The student used AI in two actions (pressing a get visualizations button in Colab and entering a chat prompt “How do I clean my dataset”), reflecting AI usage ai_suggest_steps_or_plan and ai_explain_concepts and indicating that the student formed the intent of the first data cleaning episode.
6.1.2. Semantically meaningful log analysis.
Raw logs consist of fragmented sequences of actions that lack meaningful semantics. To recover actual student usage and rationale, we analyze user logs through a hierarchy of episodes, steps, and log actions (as shown in Figure 4B).
Specifically, we segmented logs into episodes — goal-driven units of activity (e.g., ”remove duplicates in data column X”). Inspired by Norman’s Stages of Action for user cognitive processes (Tenner, 2015), we further broke each episode into four steps — forming intent, expressing the intent as an input (e.g., writing code or prompt), obtaining and understanding output (e.g. inspecting the code execution results or errors), and assessing output (e.g., deciding if the intent is sufficiently fulfilled). Each step is achieved through a sequence of captured log actions. Within each step, we annotated whether the student encountered challenges, whether the step succeeded or failed, what AI usage behavior occurred, and what additional AI usage behavior could have helped, or missed opportunities of benefiting from more effective AI usage.
Develop the codebook.
To answer “how are users using LLMs”, we identify the core AI usage behaviors by developing a codebook (Table 3). Two authors thematically analyzed a subset of homework logs and screen recordings, following a hybrid of deductive and inductive thematic analysis approach (Braun and Clarke, 2006; Fereday and Muir-Cochrane, 2006). To maximize verifiability, we used five logs and screen recordings from HW0 and five logs and think-aloud recordings from HW1-4. The AI code types in our codebook included both the AI use objectives (improving, coding, evaluating, explaining), and quality of the corresponding student prompts, as using an AI is not the same as using the AI well (echoing observations in literature (Zamfirescu-Pereira et al., 2023; Ma et al., 2025)).
Table 3. Codebook for AI usage behaviors. Code reflect different usage objectives (to have AI improve plans, assist coding, explain concepts, or evaluate deliverables), as well as usage quality (whether prompt was clear enough, etc.). For the prompt quality category, we provide negative examples since other rows already demonstrate positive examples.
Code Definition and Examples
Improve (Plan)
ai_breakdown_intent Decompose a complex goal into manageable pieces to form the correct task intent. E.g., “How to improve performance of model. give step by step.”
ai_improve_prompt Refine the wording of a prompt for clarity, specificity, and completeness before generation. E.g., “plz improve my prompt: [prompt]”
ai_suggest_steps_or_plan Provide a step-by-step plan or workflow (e.g., data cleaning $\rightarrow$ modeling $\rightarrow$ evaluation) for the task. E.g., “How to do data cleaning?” or ”How to build ML models?”
Code
ai_generate_code Produce code that implements a clearly stated action to achieve the intended output. e.g., “Write the code to calculate the distribution of the cars by location.”
ai_edit_partial_code Modify only a selected snippet or function instead of rewriting the whole cell or file. e.g., [Ask AI to only edit part of the code instead of regenerating the entire cell.]
Explain
ai_explain_bug_or_error Explain an error or traceback and outline a concrete fix. E.g., Copy-paste error message and use ”explain error” to understand and fix it.
ai_explain_code_or_api Interpret code or clarify specific APIs (e.g., describe what a function does and how to use it). E.g., “Explain what pd.get_dummies does”
ai_explain_concepts Provide plain-language explanations of data science or machine learning concepts. E.g., “How do data cleaning steps work?”
Eval
ai_critique_output Evaluate results or insights for correctness, alignment with goals, and suggest improvements. E.g., “Check my insight and provide improvements”, “Is it a good model to work with?”
Prompt (Quality)
clear_instruction_in_prompt Write prompts with explicit action verbs and desired outcomes (avoiding vague instructions like “fix it”). “Fix it” without saying what ”it” is would be unclear.
provide_context_in_prompt Include necessary details (e.g., dataframe names, column names, prior steps) and avoid vague pronouns to prevent misinterpretation. E.g., “Don’t make up non-existent column name” without providing actual context does not help resolve hallucination.
Scale up with automated log annotation
Because manual annotation was time-intensive [^3] Even with the video and with one author already familiar with the session, on average, it took ¿ two hours to go through each log. It is even more difficult to decipher logs that do not come with video recordings, due to, e.g., readability issues (certain log events would contain more notebook context information with code cell metadata that are hard to manually read through)., we scaled up by using an LLM to label log steps with the codebook (claude-sonnet-4-20250514 for its reasoning capability, with temperature=0 for deterministic generations, full prompts in Appendix D). We had the model identify the challenges as well as AI usage — either the student used AI (usedused), or the student missed an AI use opportunity (missedmissed) [^4] Our complete annotation output also contains a series of other features we used for supporting the model’s inference time reasoning procedure and the later-on human verification, e.g., log_id, the reasoning behind certain label, etc. Complete prompt with the output format and other hyperparameters are in Appendix D..
Validation of LLM annotation.
Our annotation produced, on average, 7.6 episodes per log file, with 5.6 steps per episode. Out of a total of 7315 log events, there are 1483 steps (form intent, write code, understand output, and assess output) across all episodes, and 920 (62.0%) of them are annotated with AI usage ($used+missed$). The distribution of average AI usage frequencies by all the codes is shown in Figure 4. An author verified 10% of the log annotations, and LLM achieved an accuracy of 75% (specifically, 81.3% accuracy for step success, 76.7% for encountered challenge, and 70.9% for AI use behaviors annotations).
6.2. Analysis results: AI usage change across steps and expertise
@@
Figure 5. Distribution of the AI usage (both usedused and missedmissed) aggregated by code type and episode steps.
@@
As shown in Figure 5, there is a clear relation between how AI is used and when. Behaviors associated with AI improving (e.g., ai_suggest_steps_or_plan) occur most frequently when students try to form intent. We see similar trends for AI coding during forming input, AI explaining during understand, and AI evaluating during assess step.
Looking more into the ratio of AI used vs. missed, we notice students often missed opportunities to use AI for explanation and evaluation, despite explain code and explain errors being supported by one-click features in Colab notebook. As shown in later analysis (Section 7.1), we suspect students felt overloaded by the information from Gemini.
@@
Expertise vs. AI usage: Experts use AI more strategically, and have more cognitive budget to explore.
From our logs, we found that expertise got embedded into AI usage in two ways (echoing prior research (Palta et al., 2025)):
  • Technical experts generate code with AI more successfully. As shown in Figure 6A, compared to novices (79%), experts had a 90% success rate in steps where they used AI to generate or edit code. Experts may successfully generate executable code on their first try, or are able to easily debug it. Further analyzing with student grades, it seemed like novice rely more on AI coding being accurate — novice achieving high scores got 82% AI coding success, 4% higher than novices in the low score group (split based on the student homework median grades of 62.2). In contrast, experts achieving high scores actually had slightly lower success on AI coding steps (90% vs. 92%).
  • Technical experts better describe their needs. As shown in Figure 6B, twice as many experts’ prompts were noted to include quality attributes of clear_instruction, decomposed_task, and context. Experts’ prompt was noticeably better, which could result from experts having more domain knowledge (Guo et al., 2024; Palta et al., 2025). Cross-comparing with the high vs. low performing student grade group, we confirmed that prompts noted with high quality attributes were largely contributed by experts who achieved a high score.
Figure 6. (A) Experts achieve a higher step success rate than novices (90% ¿ 79%) when using AI to assist with coding (ai_generate_code and ai_edit_partial_code); Dividing by grades shows more variations. (B) Experts with high grades contributed to most prompts that include high quality attributes.
@@
Figure 7. (A) While experts and novices used AI with similar frequencies, novices were more driven by resolving challenges with AI, and experts were more able to resolve challenges independently without using AI. (B) Compared to novices, experts were more likely to ask AI to explain only when they encountered challenges, and proactively plan and improve even without challenges. These reflect experts’ more strategic distribution of cognitive effort.
We also saw that technical experts using AI more strategically. As shown in Figure 7A, experts and novices used AI at similar frequencies (in around 36% of steps). However, experts were less pressed into using AI when facing challenges — experts’ AI usage (44%) was 10% less than novices for resolving challenges. As shown in Figure 7B, experts ask AI for explanation more often than novices when they ran into challenges (75% vs. 57%); meanwhile, their requests for AI to improve or plan were less tied to challenges than novices (43% ¡ 64%). Experts could be more aware of their goals, or they had more cognitive load to try themselves due to familiarity with the task. Indeed, experts were more capable of resolving challenges without using AI than novices (Figure 7A).

### v2

@@ 4. Method
Figure 2. Distribution of self-rated technical experience, LLM experience, and communication experience scores.
Three bar charts showing distributions of students’ self-rated technical experience, LLM experience, and communication experience. Technical and LLM experience vary widely, while communication ratings cluster on the higher end.
We used a mix of quantitative and qualitative methods to answer our research questions. RQ1 (Do LLMs close the experience gap?) was motivated by the premise that LLMs could level the field for less-experienced students, prompting us to statistically examine whether performance gaps persist and why. RQ2 (How do differences in experience translate to AI usage behaviors?) sought to understand students’ real-world AI use patterns in behavioral logs and how these relate to experience levels and performance. RQ3 (What can we train students on AI usage?) was built on these findings to identify which AI competencies require deeper pedagogical investment to truly help students benefit from AI use. Here, we provide a method overview, and some details (such as the regression factors and survey questions) are presented in the corresponding sections:
  • RQ1 (Section 5): To assess whether different experience makes a significant difference in grades, we conduct statistical testing, including non-parametric tests and regression on student homework data and experience scores;
  • RQ2 (Section 6): To characterize students’ AI usage behavior in relation to their task completion sequences, experience, and grades, we perform thematic analysis to identify key AI behavior types, use them to annotate student logs (Section 4.3), and analyze the resulting distributions;
  • RQ3 (Section 7): To further identify skills essential for effective AI usage, we qualitatively analyzed survey responses as well as the aforementioned student log behaviors, identifying skill challenges and learning opportunities across conceptual, procedural, dispositional, and metacognitive dimensions.
@@ 4.1. Participants
A total of 41 students completed HW0, with a slight attrition to 36 students by HW4. Our analysis in this work was performed on these 36 students except for log analysis (Section 4.3), in which we filtered data to ensure quality (see details about sample size and data sources for each method in Appendix B). Students came from a wide range of fields, including finance, telecommunications, journalism, physics, music, and computer science. This diversity highlights both the promise and the challenge of designing AI supports that work for “everyone.”
We used the widely adopted self-report practice and common questions in prior literature, such as “How many years of experience do you have in using Python?” (Chen et al., 2024a; Lum et al., 2025; Chen et al., 2024b) in our surveys. We collected pre-survey data on different self-reported experiences, including implementation (e.g., Python), domain (e.g., data science), toolkit (e.g., Colab/Jupyter Notebook, pandas, which are the specific environment and libraries used for our tasks), model (e.g., GenAI/LLM proficiency), and non-technical skills (e.g, communication abilities). We initially aimed to analyze whether these factors influence effective LLM use in different ways. However, in our class, some skills were moderately correlated (e.g., a Spearman correlation $\rho=0.52$ for Python and data science items, and $\rho=0.72$ between data science and toolkit items), making it difficult to disentangle their independent contributions (more on future work in Section 8.2).
Therefore, we operationalize students’ technical experience by grouping the survey answers to programming, data science, and toolkit experience questions, such as year of experience using Python and “How would you assess your level of expertise in [Exploratory data analysis]?”. Similarly, we merged highly correlated items into composite measures of LLM experience (e.g., “How would you assess your level of expertise in [GenAI/LLM]?”) and communication experience (e.g., “I am confident in my communication ability in general.”). Items that were highly skewed toward zero or not informative were excluded, and we also used methods such as principal component analysis (PCA) (Greenacre et al., 2022) and variance inflation factors (VIF) (Marcoulides and Raykov, 2019) to empirically verify our grouping (more details in Appendix D). Specifically, multicolinearity is unlikely among the technical, LLM, and communication experience variables (Figure 10b), enabling further analysis of these skills. The distribution of these experiences among students is shown in Fig. 2, where students were classified as novices or experienced using a median split of self-ratings.
Figure 3. (A) Example raw log event format. (B) Example annotated log format. The student used AI in two actions (pressing a get visualizations button in Colab and entering a chat prompt “How do I clean my dataset”), reflecting AI usage ai_suggest_steps_or_plan and ai_explain_concepts and indicating the intent of the first data cleaning episode.
Example raw log event in JSON format showing metadata, prompt text, and LLM output, alongside an annotated version labeling AI usage behaviors and the inferred intent-formation step in a data-cleaning episode.
@@ 4.2. Data Collection
We administered structured surveys at multiple points in the course (see Appendix C for full questions):
  • Pre-Survey (start of semester): Collected demographics, academic and professional backgrounds, Likert-scale questions on self-perceived experience, interest, confidence in personal vs. GenAI ability in different fields, etc.
@@
  • Final Post-Survey (end of semester): Collected self-assessed experience, interest, and confidence across domains (similar to pre-survey). Probed on similarities and differences across tasks, what they learned, and what they still wanted to learn.
@@
We deployed a custom instrumentation in Colab to capture notebook interactions (code executions/errors), selected UI events (pressing explain-code/error buttons), and LLM use (prompts/responses). Students installed a browser script that intercepted requests to Colab’s converse and generateCode endpoints and captured clickstream events. The script added a “Download History” button to the Colab toolbar and stored code cells, chat messages, event metadata, and responses to browser localStorage, which students exported as JSON files upon submission (or auto-save when the storage is full). In total, we collected 16,315 logged events (e.g., Fig. 3A) and 44 hours of screen recordings across five assignments. Logs captured a rich sequence of AI interactions, code edits, and executions, supporting both quantitative distributional analyses and qualitative episode-based coding (Section 4.3).
4.3. LLM-assisted Log Data Annotation
Table 2. Codebook for AI usage behaviors. Code reflects different conceptual (what to use AI for) and procedural (how to use AI effectively through prompting) knowledge dimensions. For the prompt (quality) category, we provide negative examples since other rows already demonstrate positive examples.
Dim. Category Code Definition and Examples
Conceptual Knowledge Dimension Improve (Plan) ai_breakdown_intent Decompose a complex goal into smaller pieces. E.g., “How to improve performance of model. give step by step.”
ai_improve_prompt Refine the wording of a prompt for clarity and specificity. E.g., “plz improve my prompt: [prompt]”
ai_suggest_steps_or_plan Provide a step-by-step workflow (e.g., data cleaning → modeling). E.g., “How to do data cleaning?” or “How to build ML models?”
Code ai_generate_code Produce code that implements a requested action. E.g., “Write code to calculate the distribution of cars by location.”
ai_edit_partial_code Edit a specific snippet or function rather than rewriting the whole cell.
Explain ai_explain_bug_or_error Explain an error or traceback and outline a concrete fix. E.g., Copy-paste error message and use ”explain error” to understand and fix it.
ai_explain_code_or_api Interpret code or explain APIs, such as describing what a function does and how to use it. E.g., “Explain what pd.get_dummies does.”
ai_explain_concepts Provide explanations of concepts. E.g., “How do data cleaning steps work?”
Eval ai_critique_output Evaluate results or insights for correctness, alignment with goals, and suggest improvements. E.g., “Check my insight and provide improvements”, “Is it a good model to work with?”
Procedural Prompt (Quality) clear_instruction_in_prompt Write prompts with explicit actions and desired outcomes. Example of a poor prompt: “Fix it” without identifying what “it” is.
provide_context_in_prompt Include necessary details (e.g., dataframe names, prior steps) and avoid vague pronouns. E.g., “Don’t make up non-existent column name” without providing actual context does not help resolve hallucination.
Codebook listing AI usage behaviors across conceptual and procedural knowledge dimensions with definitions and examples, including planning, coding, explanation, evaluation, and prompt-quality categories.
To answer how students use LLMs (RQ2) and what we should teach on effective LLM usage (RQ3), we need to identify the core AI usage behaviors. To this end, we analyze students’ log data to develop a codebook to capture behavioral patterns in authentic problem-solving processes.
Log Parsing
Raw logs consist of fragmented sequences of actions that lack meaningful semantics. To recover actual student usage and rationale, we analyze user logs through a hierarchy of episodes, steps, and log actions (as shown in Fig. 3B). Specifically, we segmented logs into episodes — goal-driven units of activity (e.g., ”remove duplicates in data column X”). Inspired by Norman’s Stages of Action for user cognitive processes (Tenner, 2015), we further broke each episode into four steps — forming intent, expressing the intent as an input (e.g., writing code or prompt), obtaining and understanding output (e.g. inspecting the code execution results or errors), and assessing output (e.g., deciding if the intent is sufficiently fulfilled). Each step contains a sequence of log actions. Within each step, we annotated whether the student encountered challenges, whether the step succeeded or failed, what AI usage behavior occurred, and what additional AI usage behavior could have helped, or missed opportunities of benefiting from AI usage.
Codebook for AI usage behavior analysis.
To develop the codebook (Table 2), two authors thematically analyzed a subset of homework logs and screen recordings, following a hybrid of deductive and inductive thematic analysis approach (Braun and Clarke, 2006; Fereday and Muir-Cochrane, 2006). To maximize verifiability, we used five logs and screen recordings from HW0 and five logs and think-aloud recordings from HW1-4. The AI code types in our codebook are characterized by two dimensions: conceptual knowledge on what to use AI for (using AI to improve plans, code, explain concepts, or evaluate deliverables), and procedural knowledge of how to use AI via good quality prompts, as using an AI is not the same as using the AI well.
Our codebook integrates and extends prior work in multiple layers — it links human-AI collaboration with Bloom’s knowledge dimensions (Krathwohl, 2002), which enables pedagogical alignment (more in Section 7). The knowledge dimensions offer a complementary perspective to the current prompt coding schema based on Bloom’s cognitive process (Chen et al., 2024b). The code categories build on and decompose existing themes that characterize human-AI collaborative behaviors, including planning, evaluating, and coding (Chen et al., 2024a). Our prompting codes also echo existing strategies such as “be specific,” “write accurate requirement” (clear_instruction_in_prompt) and “set the scene” (provide_context_in_prompt) (Tassoti, 2024; Ma et al., 2025).
Scale up with automated log annotation
Because manual annotation was time-intensive [^2] Even with the video and with one author already familiar with the session, on average, it took ¿ two hours to go through each log. It is even more difficult to decipher logs that do not come with video recordings, due to, e.g., readability issues (certain log events would contain more notebook context information with code cell metadata that are hard to manually read through)., we scaled up by using an LLM to label log steps with the codebook (claude-sonnet-4-20250514 for its reasoning capability, with temperature=0 for deterministic generations, full prompts in Appendix E). We had the model identify the challenges as well as AI usage — either the student used AI (usedused), or the student missed an AI use opportunity (missedmissed) [^3] Annotation output contains other features to support the model’s inference time reasoning and human verification, e.g., id, reasoning behind certain label, etc. Complete prompt with the output format and hyperparameters are in Appendix E..
Data filtering.
To ensure the reliability of our quantitative log analyses that examine experience effects and pre-post instructional changes on interaction behaviors, we applied a series of filters to improve data quality. Specifically, we filtered out [^4] We required students to use the built-in Gemini and avoid using other LLMs and added the precaution of asking students to self-report their additional LLM model use, if any. However, we still noticed students using external AI tools, which tend to result in reasonable grades but no or few user logs captured, or logs with abnormal behaviors, e.g., copy-pasted large blocks of code with unfounded originality. students who self-reported other tool use like ChatGPT or displayed other tool use in recording (if available), as well as those who disengaged or spent too much time working on tasks without AI, reflected as log length being less than the median. To enable comparisons of behavioral changes before and after instructions, we only apply the quantitative analysis over logs on T1 data cleaning and T2 EDA, as most interactions come from tasks 1 and 2 in HW0 due to time limits. As a result, we only annotated 7,315 logged events from 28 students across HW0-2.555Each homework may include a different subset of student submissions after filtering. We conduct log annotations over a total of 36 submissions (20 from experienced, 16 from novices). Note that “experienced” vs. “novice” denote technical experience here and for the rest of the paper.
Validation of LLM annotation.
Our annotation produced, on average, 7.6 episodes per log file, with 5.6 steps per episode. Out of a total of 7315 log events, there are 1483 steps (form intent, write code, understand output, and assess output) across all episodes, and 920 (62.0%) of them are annotated with AI usage ($used+missed$). The distribution of average AI usage frequencies by all the codes is shown in Fig. 9. An author verified 10% of the log annotations, and LLM achieved an accuracy of 75% (specifically, 81.3% accuracy for step success, 76.7% for encountered challenge, and 70.9% for AI use behaviors annotations).
5. RQ1: Do LLMs close experience gap?
Students engaged substantially with AI during coursework: 15% of logged events were prompt-related, and homework post-surveys all confirmed that students spent non-trivial amounts of time with AI tools. Despite substantial AI usage, we found that technically experienced students seem to benefit more from LLM use than novices if given ample time.
We hypothesized that students with higher technical experience would perform better in our programmatic data science tasks. To examine if using LLM helps bridge this performance gap, we compared performance between more technically experienced students and novices in the timed HW0 under conditions with and without LLMs (Table 1). A non-parametric Mann–Whitney U test (as data violated normality assumptions) (McKnight and Najab, 2010) showed a significant technical experience gap when students did not use LLMs ($p=0.027<0.05$), but the gap disappeared when LLM was used (p>0.05p>0.05) [^6] We also verified that the communication and LLM experience do not lead to any significant performance gap with or without LLM in HW0.. This suggests that under time pressure, LLMs may offset technical experience’s advantages.
@@ 5. RQ1: Do LLMs close experience gap?
Table 3. Linear mixed-effects regression predicting homework performance from different experiences. Participant and homework were used as random intercepts.
@@
Regression results table showing that technical experience significantly predicts homework grades, whereas LLM experience and communication skills do not.
To further probe the interplay between different experiences’ influence over LLM use in more authentic scenarios without the time constraints, we assess whether different experiences make a significant difference in grades in HW1–HW4.
@@
We fit a linear mixed-effects regression model (Bates et al., 2015) with grades (normalized scores) as the outcome and technical, LLM, and communication experience scores as predictors, including random intercepts for participant and assignment. As shown in (Table 3), only technical experience significantly predicted higher grades ($\beta=6.09,p=.041$). LLM and communication experience had no significant effects and did not predict grades. This suggests that even with equal access to the Colab Gemini assistant, students with stronger programming and data science backgrounds continued to outperform peers when given ample time. In other words, the technical experience gap may not be closed by an LLM in a more realistic setup, which is consistent with some prior literature (Chen et al., 2024a).
@@
Note that students in our class had a median of 4/5 for their communication experience (see Fig. 2), which means that the communication factor may not have a lot of variance to predict grades. Meanwhile, students’ median of LLM experience is 3.67/5, but everyday use of LLMs or their self-rated LLM knowledge may not straightforwardly transfer to complex programming data science tasks. Our HW0 tasks without LLM use may serve as a proxy of pre-assessment over technical experience, as partially validated by the significant performance difference between more technically experienced and novice students. Nevertheless, future work should explore more rigorous ways to measure experience and recruit a more diverse set of students (more in Section 8.2).
6. RQ2: How do experience differences translate to AI usage behaviors?
@@
The analysis in RQ1 showed that students’ self-perceived technical experience continued to predict performance even when all students had access to the Colab Gemini assistant. Here, we dive into student log data to examine why: how students use AI during workflows, and how it relates to different levels of technical experience?
@@
Figure 4. Distribution of the AI usage (both usedused and missedmissed) aggregated by code type and episode steps. Bar length represents the number of steps involving this AI use, while the percentages denote the proportion of usedused vs. missedmissed opportunities. 100% means that all AI usage is usedused (if blue) or missedmissed opportunities (if grey).
Stacked bar chart showing AI usage across four episode stages—intent, input, understand, assess—with bars indicating used versus missed opportunities for planning, coding, explanation, evaluation, and prompt-quality behaviors.
@@
There is a clear relation between how AI is used and when (Fig. 4). Behaviors associated with improving (e.g., ai_suggest_steps_or_plan) occur most frequently when students try to form intent. We see similar trends for AI coding during input, AI explaining during understand, and AI evaluating during assess step.
Looking more into the ratio of AI usage, we notice students often missed opportunities to use AI for explanation and evaluation, despite explain code and explain errors being supported by one-click features in Colab notebook. As discussed in Section 7, we suspect students felt overloaded by AI output.
@@
Experience vs. AI usage: More technically experienced students could use AI more strategically, and have more cognitive budget to explore.
@@
Figure 5. (A) While experienced students and novices used AI with similar frequencies, experienced students might be more able to resolve challenges without using AI. (B) Experienced students might strategically distribute effort to ask AI to plan and improve or explain, different from novices.
Two sets of bar charts comparing novices and experienced students. One shows that experienced students resolve more challenges without AI; the other shows differences in how each group requests planning, explanation, and coding support from AI.
Figure 6. (A) Experienced students achieve a higher step success rate than novices when using AI to assist with coding (ai_generate_code and ai_edit_partial_code); Dividing by grades shows more variations. (B) Experienced students with high grades contributed to most prompts with high prompt (quality) attributes. High or low grades are split based on the median.
Bar charts showing experienced students’ higher success rates in AI-assisted code generation and more frequent use of high-quality prompt attributes compared to novices and lower-performing students.
As shown in Fig. 5A, experienced students and novices used AI at similar frequencies (in around 36% of steps). However, experienced students might be less pressed into using AI when facing challenges — they were more capable of resolving challenges without using AI than novices (Fig. 5A). Nevertheless, when experienced students did use AI (Fig. 5B), they might ask AI for explanation more often than novices when they ran into challenges (75% ¿ 57%). Meanwhile, their requests for AI to improve or plan seemed less tied to challenges than novices (43% ¡ 64%). This different AI use strategy may reflect technically experienced students’ clearer awareness of when and what type of help is needed. They may be better able to identify the challenges and seek targeted clarification and better request planning support proactively before encountering challenges, compared to novices.
Moreover, our logs reveal patterns echoing prior research (Guo et al., 2024; Palta et al., 2025), such as experts are better at generating code using AI and prompting AI, which could result from having more domain knowledge. As shown in Fig. 6A, technically experienced students generate code with AI more successfully (90% ¿ 79%). Further analyzing with student grades, it seemed like novices might rely more on AI coding being accurate to achieve higher scores — novices with higher grades also had a higher AI coding success rate than novices with lower scores (82%¿ 78%). In contrast, experienced students achieving high scores actually had slightly lower success on AI coding steps (90% ¡ 92%). As shown in Fig. 6B, technically experienced students better describe their needs. Prompts from experienced students who achieved high scores included twice as many quality attributes of clear_instruction, decomposed_task, and provide_context.